Mary Jarvis is a single individual who is working on filing her tax return for the previous year. She has assembled the following relevant information:

She received $78,000 in salary.

She received $14,500 of dividend income.

She received $4,200 of interest income on Home Depot bonds.

She received $25,000 from the sale of Disney stock that was purchased 2 years prior to the sale at a cost of $5,300.

She received $11,500 from the sale of Google stock that was purchased 6 months prior to the sale at a cost of $5,400.

Mary receives one exemption ($4,000), and she has allowable itemized deductions of $7,500. These amounts will be deducted from her gross income to determine her taxable income.

  • What is Mary's federal tax liability? Round your answer to the nearest cent. Do not round intermediate calculations.

  • What is her marginal tax rate? Round your answer to 1 decimal place.

  • What is her average tax rate? Round your answer to 2 decimal places.
Answer #1

Part A

Tax liability calculation

Amount ($)
Salary 78,000
Dividend Income 14,500
Interest Income 4,200
Profit from sale of Disney Stock ( 25000-5300 ) 19,700
Profit from sale of Google Stock ( 11500-5400 ) 6,100
Less : Deductions
Exemption (4,000)
Itemized deduction (7,500)
Tax Liability 111,000

Part B

Marginal Tax Rate

Assume below mentioned is tax rate ( because tax rate is not given in question )

Tax Table for Single Individuals:
Taxable Income Amount Paid on Base Percentage on Excess over Base
$0.00 $0.00 10.00%
$9,225.00 $922.50 15.00%
$37,450.00 $5,156.25 25.00%
$90,750.00 $18,481.25 28.00%
$189,750.00 $46,075.25 33.00%
$411,500.00 $119,401.25 35.00%
$413,200.00 $119,996.25 39.60%

The taxable income of Mary calls in the bracket of $90750- 189750

Therefore Marginal Tax Rate is 28%

Part C

Average Tax Rate = Tax payable / Taxable Income

Tax payable = $18481.25+ ($111000-$90750) ×28% = 24,151

Therefore, Average Tax Rate = 24,151 / 111,000 = 21.75%

Average Tax Rate is 21.75%